how many grams of ammonium nitrate must be weighed out to make 451 grams of a 58.0% solution of water?
Answer #1

Total weight of solution = 451 gm

Ammonium nitrate weight % solution = 58 %

Water weight % = 100 - 58 = 42 %

Mass of Ammonium nitrate = 58 x 451 / 100 = 261.58 gm

Mass of water = 42 x 451 / 100 = 189.42 gm

Hence the solution can be made with 261.58 gm of ammonium nitrate and 189.42 gm of water.